OpenStudy (love_to_love_you):

Identify the points at which f(x) is differentiable. f(x) = |x + 3|; (-6, 3)

OpenStudy (gluttrell):

what grade are u in.

OpenStudy (love_to_love_you):

10th

OpenStudy (gluttrell):

same what math are u taking

OpenStudy (love_to_love_you):

Pre-cal

OpenStudy (gluttrell):

oh wow u are really smart.

OpenStudy (love_to_love_you):

lol not really

OpenStudy (gluttrell):

yeah im in geometry.

OpenStudy (love_to_love_you):

ah cool

OpenStudy (love_to_love_you):

@confluxepic can you help?

OpenStudy (jdoe0001):

hmm what's the (-6,3? anyway?

OpenStudy (gluttrell):

see if this will help u https://www.sagemath.org/calctut/differentiability.html

OpenStudy (jdoe0001):

an absolute value graph, will have an "abrupt shfit or turn", namely a "cusp" at the vertex so. is the only place it's not differentiable at

OpenStudy (love_to_love_you):

. ?

OpenStudy (jdoe0001):

\(\huge ?\)

OpenStudy (love_to_love_you):

so if I graph it the shift is the answer?

OpenStudy (gluttrell):

ok if that did not help u ignore me

OpenStudy (jdoe0001):

ahemm the "cusps" or the vertex in this case is the only place is NOT differentiable is differentiable everywhere else

OpenStudy (love_to_love_you):

It's loading @gluttrell

OpenStudy (love_to_love_you):

ohh, sorry I read that wrong

OpenStudy (love_to_love_you):

So it's differentiable everywhere except at x = -3?

OpenStudy (jdoe0001):

bear in mind differentiable -> has a smooth line over which a tangent line can easily sit a cusp or an abrupt edge or break, is not smooth, and thus you can get a tangent line there, and thus is not differentiable you can pick those up by simply setting the denominator of the derivative to 0 and solve for "x"

OpenStudy (gluttrell):

sort of ur on the right track.

OpenStudy (jdoe0001):

yeap, -3, 0 is the vertex, so, is not differentiable at x = -3 everywhere else, is

OpenStudy (love_to_love_you):

Thank you, the both of you

OpenStudy (jdoe0001):

and thus you can't get a tangent line there, and thus is not differentiable correcting typo, "can't" I meant :)

OpenStudy (jdoe0001):

yw

OpenStudy (sshayer):

\[\frac{ d }{ dx }\left| f(x) \right|=\frac{ f \left( x \right) }{ \left| f \left( x \right) \right|}f \prime \left( x \right)\] \[\frac{ d }{ dx }\left| x+3 \right|=\frac{ x+3 }{ \left| x+3 \right| }*1=\frac{ x+3 }{ \left| x+3 \right| }\] it is not differential where denominator |x+3|=0 or x=-3 derivative does not exist at x=-3

ramen:

did you just have a conversation with youself?

justjm:

The better way to answer this question is to convert the abs. val. function into a piece wise and evaluate continuity and differentiability using the definition of continuity for the function f(x) and function f ' (x). Also abs. value functions are never differentiable at the vertex because although it is continuous, the slope function has different right and left hand limits.